OpenStudy (jotopia34):

Use a double integral to find the area of the region. The region inside the circle (x − 5)^2 + y^2 = 25 and outside the circle x^2 + y^2 = 25 Which one is the inside Integral and how do I find the bounds for it. Is x=r, or is x=rcos(theta)? I'm confused.

OpenStudy (jotopia34):

\[(x-5)^2+y^2=25... and...x^2+y^2=25\]

OpenStudy (anonymous):

inside the circle, r <= 5

OpenStudy (anonymous):

remember, x = r cos theta and y = r sin theta

OpenStudy (jotopia34):

okay. But is one of the limits of integration 0 and 5?

OpenStudy (anonymous):

yeah, for inside the circle. 0 <= r <= 5

OpenStudy (jotopia34):

which equation is the "inside" intagral? For example, to keep dbl integrals straight I write them \[\int\limits_{a}^{b}[\int\limits_{c}^{d}f(x,y) say, dx]dy\]. Here, the inside integral is the dx.

OpenStudy (anonymous):

Rewrite your function in terms of r and theta... and dx dy becomes r dr dtheta \[\left( x-5\right)^2 +\left( y \right)^2= 5^2 \Rightarrow x-5 = r \cos \theta \text{, }y=r \sin \theta\]

OpenStudy (jotopia34):

I did that and found r=10cos(theta)

OpenStudy (anonymous):

?

OpenStudy (jotopia34):

sorry, i changed x2+y2=25 into P.C and got r=+-5, rejectiing -5, so some boundary is 0,5. Not correct?

OpenStudy (jotopia34):

what is getting me, is that there are 2 equations. One that's the inside of the circle and one that's the outside. Which one is getting integrated on my "inside" integral

OpenStudy (anonymous):

same equation... different limits of integration

OpenStudy (jotopia34):

Again, which one is inside and why?

OpenStudy (jotopia34):

:( I'm in deep doodoo.

OpenStudy (anonymous):

if you want the area of the circle, you just use 1 as the function. \[\int\limits_{0}^{2 \pi}\int\limits_{0}^{5}1r\,dr\,d \theta = \cdots\]

OpenStudy (anonymous):

that's for the area inside the circle. for the area outside the circle, the limits of integration on r would be from 5 to infinity.

OpenStudy (jotopia34):

It says find the area inside the circle AND the area outside. so do I still just use equn 1? iF SO, how do I know not to use x^2+y^2=25 as my Integrand?

OpenStudy (anonymous):

when using a double integral merely to find an area, you integrate over the appropriate region using only dA as the integrand...